Circular No. 8616

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005ex–2005gj AND 2005go–2005gy CBETs 247 and 254 contains discovery information (astrometry, pho-

tometry, redshifts, spectroscopic classifications) for 50 new supernovae discovered in September and early October in the course of the Sloan Digital Sky Survey II and communicated by J. Frieman (Fermilab and University of Chicago). All the objects are now apparently fainter than red mag 19 except 2005gj, a probable type-IIn supernova that may currently be near mag 18. SN 2005fk is a type-Ib/c hypernova, SN 2005gi is a type-II supernova, and the rest apparently are of type Ia.

SUPERNOVAE 2005gm, 2005gn, 2005gz, 2005ha, 2005hb Five apparent supernovae have been discovered on unfiltered CCD

frames: 2005gm by P. Luckas, O. Trondal, and M. Schwartz (cf. IAUC 8575) with the 0.35-m telescope at Perth, and the rest by R. R. Prasad, K. Shimasaki, E. Lee, and W. Li (LOSS/KAIT; cf. IAUC 8611).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005gm Oct. 9.86 3^{h}42^{m}39.^{s}30 − 6^{o}23′04′′.0 18.0 12′′.5 W, 8′′.9 S
2005gn Oct. 12.52 5 48 49.07 −24 22 45.5 18.4 7′′.8 E, 0′′.6 S
2005gz Oct. 11.18 21 05 20.69 − 7 50 47.5 18.3 2′′.9 S
2005ha Oct. 13.52 6 21 49.04 + 0 21 55.6 18.3 21′′.2 W, 2′′.2 S
2005hb Oct. 13.53 6 45 31.64 +25 49 58.0 19.5 18′′.8 E, 7′′.6 N

Approximate unfiltered magnitudes from the respective groups: SN 2005gm in NGC 1423, 2004 Oct. 16.31 UT, [19.0 (Schwartz, 0.81-m telescope at Patagonia, AZ); 2005 Oct. 12.38, 18.0. SN 2005gn in ESO 488-G30, 2004 Mar. 10.15, [19.0; 2005 Sept. 12.55, [18.0; Oct. 13.53, 18.0. SN 2005gz in MCG –01-53-22, Sept. 23.19, [20.0; Oct. 2.20, [17.5; 14.18, 18.0. SN 2005ha in UGC 3457, Oct. 3.51, [19.0, 14.50, 18.0. SN 2005hb in UGC 3531, 2004 Mar. 12.28, [20.0; 2005 Sept. 28.53, [19.5; Oct. 3.54, [18.8; 14.50, 19.2.

COMET C/2005 Q6 (SOHO) Further to IAUC 8615, K. Battams reports measurements for a very

faint comet (with a slightly irregular shape) that does not belong to any known group, found on SOHO website images.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 Q6 Aug. 19.121 9^{h}45.^{m}4 +11^{o}43′ C2 HS 2005-T55
2005 October 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green